OpenStudy (anonymous):

what fraction of 2 1/3 is 2 1/9? a)19/21 b)18/22 c)17/19 d)23/37

OpenStudy (anonymous):

2 1/3 = 7/3 2 1/9 = 19/9 The question can be rephrased as \[ \frac{7}{3} x = \frac{19}{9} \] solve for x you get x = (3*19)/(7*9)=19/21 check out http://www.oojih.com for free study guide and free review questions